Due to the current gap in continued funding from the U.S. National Science Foundation (NSF), the NSF Unidata Program Center has temporarily paused most operations. See NSF Unidata Pause in Most Operations for details.
NOTE: The nws-changes
mailing list is no longer active. The list archives are made available for historical reasons.
Product Outage/Anomaly: FRAC MetOp-A, Issued: September 17, 2010 0637 UTC.eml
Subject: Product Outage/Anomaly: FRAC MetOp-A, Issued: September 17, 2010 0637 UTC From: ESPCOperations <ESPCoperations@xxxxxxxx> Date: Fri, 17 Sep 2010 02:35:38 -0400 To: _NESDIS OSDPD ESPC OUTAGE Notification <ESPC.Notification@xxxxxxxx> *Topic: *MetOp-A (FRAC) B2029192.SV (Update) *Date/Time**: *September 17, 2010 0637 UTC* * *Product(s) or Data Impacted: * A-DCS/2 AMSU-A ASCAT AVHRR GRAS GOME HIRS IASI MHS S&R SEM **** *Date/Time of Initial Impact: *September 17, 2010 0000 UTC* * *Date/Time of Expected End: *Unknown* * *Length of Outage: *Until Further Notice* **Details/Specifics of Change: *Network Operations Center (NOC) has informed ESPC of an outage due to the data router server being down. They are currently investigating the problem.*
**Update#1: *ESPC is still not receiving MetOp-A (FRAC) data from Network Operations Center (NOC).*
**Contact Information for Further Information:* ESPC Operations at ESPCOperations@xxxxxxxx at 301-817-3880 * *
*Web Site(s) for applicable information:* Enter applicable web site address/es as necessary
Product Outage/Anomaly: FRAC MetOp-A, Issued: September 17, 2010 0637 UTC.eml
Subject:Product Outage/Anomaly: FRAC MetOp-A/ Blind GAC, Issued: September 17, 2010 0837 UTC (Update) Correction of initial impact time
From: ESPCOperations <ESPCoperations@xxxxxxxx> Date: Fri, 17 Sep 2010 04:40:29 -0400 To: _NESDIS OSDPD ESPC OUTAGE Notification <ESPC.Notification@xxxxxxxx> *Topic: *MetOp-A (FRAC) B2029192.SV (Update) *Topic: ** *MetOp-A (FRAC) and Blind GAC (Update) *Date/Time**: *September 17, 2010 0837 UTC* * *Product(s) or Data Impacted: * A-DCS/2 AMSU-A ASCAT AVHRR GRAS GOME HIRS IASI MHS S&R SEM *****Date/Time of Initial Impact: *September 17, 2010 0400 UTC (Time Correction from 0000 UTC to 0400 UTC)* *
*Date/Time of Expected End: *Unknown* * *Length of Outage: *Until Further Notice* **Details/Specifics of Change: *Network Operations Center (NOC) has informed ESPC of an outage due to the data router server being down. They are currently investigating the problem.*
**Update#1: *ESPC is still not receiving MetOp-A (FRAC) data from Network Operations Center (NOC).
*Update#2:* ESPC is still not receiving MetOp-A (FRAC) and Blind GAC data. NOC has informed ESPC that AT & T is still working on the problem. * *
*Contact Information for Further Information:* ESPC Operations at ESPCOperations@xxxxxxxx at 301-817-3880 * *
*Web Site(s) for applicable information:* Enter applicable web site address/es as necessary
Product Outage/Anomaly: FRAC MetOp-A, Issued: September 17, 2010 0637 UTC.eml
Subject:Product Outage/Anomaly: FRAC MetOp-A/ Blind GAC, Issued: September 17, 2010 0930 UTC (Update#3 )
From: ESPCOperations <ESPCoperations@xxxxxxxx> Date: Fri, 17 Sep 2010 05:33:28 -0400 To: _NESDIS OSDPD ESPC OUTAGE Notification <ESPC.Notification@xxxxxxxx> *Topic: *MetOp-A (FRAC) B2029192.SV (Update)*Topic: ** *MetOp-A (FRAC) and Blind GAC (Update#2). Blind GAC is now been received (Update#3).
*Date/Time**: *September 17, 2010 0930 UTC* * *Product(s) or Data Impacted: * A-DCS/2 AMSU-A ASCAT AVHRR GRAS GOME HIRS IASI MHS S&R SEM *****Date/Time of Initial Impact: *September 17, 2010 0400 UTC (Time Correction from 0000 UTC to 0400 UTC)* *
*Date/Time of Expected End: *Unknown* * *Length of Outage: *Until Further Notice* **Details/Specifics of Change: *Network Operations Center (NOC) has informed ESPC of an outage due to the data router server being down. They are currently investigating the problem.*
**Update#1: *ESPC is still not receiving MetOp-A (FRAC) data from Network Operations Center (NOC).
*Update#2:* ESPC is still not receiving MetOp-A (FRAC) and Blind GAC data. NOC has informed ESPC that AT & T is still working on the problem.
*Update#3: *ESPC* *has received the Blind GAC for orbit B0828788 that was schedule at 0449z.
*Contact Information for Further Information:* ESPC Operations at ESPCOperations@xxxxxxxx at 301-817-3880 * *
*Web Site(s) for applicable information:* Enter applicable web site address/es as necessary
nws-changes
archives: